2.1 The per-phase parameters for a 60 Hz, 200 km long transmission line are
1
Expert's answer
2021-08-16T02:20:47-0400

The total resistance "= 2 \pi fL= 2\pi*60*310.8= 117.16 \Omega\\"

Capacitance resistance "X_c= \frac{1}{jwC}=- \frac{j}{2 \pi fC}=- \frac{j}{2 \pi *60*1.477 \mu F}"

Therefore admittance "Y= \frac{1}{X_c}=2 \pi fC=2 \pi *60*1.477*10^{-6}= 0.0005565"